OpenStudy (anonymous):

find the equation of a line with a slope of -1 which is tangent to the circle x^2 + y^2 = 9

OpenStudy (anonymous):

yohcandoit.

hero (hero):

Find the radius of the circle first, then the point where the line hits the circle. then you'll have the point and a slope, then you can find the equation of the line.

OpenStudy (anonymous):

the radius would be 3. but how would you find where the line hits the circle?

hero (hero):

The radius is the distance. You can use the distance formula to help you find it

hero (hero):

You already have the center point, right?

OpenStudy (anonymous):

the distance from the origin :O how would that help? the center point must be (0,0)

hero (hero):

In your distance formula you have d = 3, x = 0, y = 0, x_1 = 3, the only thing left to find is y_2

OpenStudy (anonymous):

huh? I'm not sure I remember the distance formula.

hero (hero):

\[d = \sqrt{(x_2 - x_1)^2+(y_2-y_1)^2}\]

OpenStudy (anonymous):

but then y = 0.. I don't see how this helps/

hero (hero):

y doesn't equal zero

OpenStudy (anonymous):

I ended up with 3 = root (0-3)^2 +(y-0)^ so -3+y = 3

hero (hero):

y = 6

OpenStudy (anonymous):

oh yeah huh...

OpenStudy (anonymous):

so x^2 + 6^2 = 9.. how am I closer to finding the line?

OpenStudy (anonymous):

hoh crap. I got a gym appointment in 15. thanks for your help, but I still don't get it.

hero (hero):

Okay

OpenStudy (anonymous):

Hey mybrainvsme please check you question again and see if they give a point where its tangent to on the circle

hero (hero):

No nayeaddo, you have to find the point

OpenStudy (anonymous):

Well you have to find the equation of the line and all you need is the y-intercept because you already have the slope. I'm not sure if I understand your method of solving for that point

hero (hero):

I'll show you

OpenStudy (anonymous):

okay

hero (hero):

\[x^2 + y^2= \sqrt{(x_2 - x_1)^2+(y_2-y_1)^2}\] Kinda like systems of equations we have to set the circle equal to the distance formula. That how you start it off. Or rather that is one approach

hero (hero):

we can replace the x^2 + y^2 with 9 and x_1, y_1 with 0

hero (hero):

The point is 3,0

hero (hero):

I found it without even doing any calculations

OpenStudy (anonymous):

(3,0) ?

hero (hero):

Yes

OpenStudy (anonymous):

So its tangent to the point (3,0) ?

hero (hero):

as far as I know

hero (hero):

That's the only point that seems to work

OpenStudy (anonymous):

yes but that doesn't really hold true then the line has a undefined slope (3,0) is an x-intercept so if the line should be tangent to that point on the circle then its a straight line

OpenStudy (anonymous):

I'm thinking that maybe calculus is involved here by taking implicit differentiation of function...I wanted to ask the questioner what class this question came from...did you get what I said before?

hero (hero):

Yeah, I'm just an amateur, but I don't think I was totally off

hero (hero):

The point is definitely not 3,0 I realize, however

OpenStudy (anonymous):

yeah...the question itself seems to be a bit vague most of the tangent line problems you're giving a specific point where line touches function..hmmm

hero (hero):

Okay, I have it

hero (hero):

It's easy, but it's not easy to explain on here.

OpenStudy (anonymous):

Here's my solution because a circle is not a function and in our case its symmetrical across both the x and y axis so I split them up into two equations: \[y = \sqrt{9 - x ^{2}} ; y = -\sqrt{9 -x ^{2}}\]

hero (hero):

Okay, you first

OpenStudy (anonymous):

You take the first derivative of the first function and set it equal to the slope of the slope of the line which is -1.

hero (hero):

Yeah, that's good

hero (hero):

There's another way to solve it without using calculus

OpenStudy (anonymous):

The tangents form triangles with the with the x-axis with short sides 3 so hypotenuse is sqrt(18) giving plus or minus 3 sqrt 2 So lines are y = -x +- 3sqrt2 http://www.wolframalpha.com/input/?i=plot+x^2+%2B+y^2+%3D+9%2Cy%3D-x-3sqrt2%2Cy%3D-x%2B3sqrt2+%2Cx%3D0%2Cy%3D0%2Cy%3Dx%2C+x%3D-5..5%2Cy%3D-5..5

hero (hero):

estudier has it

OpenStudy (anonymous):

you get\[-x / \sqrt{9 - x ^{2}} = -1\] when you solve for x you get sqrt(4.5) and y is also sqrt(4.5) okay whats your way

hero (hero):

estudier already posted it

hero (hero):

But that's not even the way I did it

OpenStudy (anonymous):

Geometry is king....

OpenStudy (anonymous):

yeah but calculus makes it funner lol I got the same answer 3sqrt(2) = 2sqrt(4.5) for the intercept

hero (hero):

The slope is negative one, the radius is three. the line from the center of the circle to the point is perpendicular to the slope of the line

OpenStudy (anonymous):

I don't I like using calculus

hero (hero):

so the equation perpendicular to the equation of the line is y = x

hero (hero):

substitute x into the equation of the circle, replacing y, then solve for x

hero (hero):

you get x = 3sqrt2/2

hero (hero):

And then you can continue solving from there

OpenStudy (anonymous):

yeah good approach too because y = x is a one-to-one function so its convenient to plug in x into equation of circle

hero (hero):

:o), I'm glad you understood what I was attempting to explain

OpenStudy (anonymous):

yes, have you studied calculus?

hero (hero):

Yes, I have, but I'm rusty

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

As u can see, not required...

hero (hero):

Precisely

hero (hero):

If I can solve it without using calculus, I do

hero (hero):

Or at least try to

OpenStudy (anonymous):

yeah, the only reason I brought it up is I wanted to find out which class the questioner was taking or from what context he posed the question because they teach tangent lines in calculus I so my approach would have been correct in that case...but I also understand what you guys are saying

OpenStudy (anonymous):

yeah..sometimes you forget the basics

OpenStudy (anonymous):

they are supposed to teach tangent lines in geometry...

hero (hero):

They're supposed to teach a lot of things in geometry, but they skip a lot of things

hero (hero):

My high school sucks

OpenStudy (anonymous):

yes but they don't introduce concept of derivative = slope in geometry I don't want to argue with you about this my point is simple and clear

OpenStudy (anonymous):

You can have complex functions that you will have to find tangent lines to as well

OpenStudy (anonymous):

Why would u want to do that?

OpenStudy (anonymous):

I said I don't want to..please can you read...the question is solved so can we stop putting unnecessary notifications for the questioner?

OpenStudy (anonymous):

Oooh, the questioner can say that themselves, no?

OpenStudy (anonymous):

lol estudier why are you so against my method...please explain..I told you understand that its not useful only if the instructor did not inform them to use a calculus approach..that's all I'm saying

OpenStudy (anonymous):

I am not against your method, if that is what floats your boat, go ahead. All I'm saying is that it's a 1 liner, no calculus required.

hero (hero):

If I were the instructor, I wouldn't care what method you use. Long as you get the correct answer and show your steps. I'd give extra credit to those who solve without calculus, lol

OpenStudy (anonymous):

haha...good for you Hero

OpenStudy (anonymous):

Nayeaddo. I'm taking precalc. This is actually part of my summer math from alg 2 lol

hero (hero):

lol

OpenStudy (anonymous):

and I still don't get it btw.

hero (hero):

I can show you on twiddla

OpenStudy (anonymous):

what's twiddla?

hero (hero):

My way is the simplest

OpenStudy (anonymous):

Just look at the link I posted, it's obvious...

OpenStudy (anonymous):

One liner.

OpenStudy (anonymous):

None of the lines on your link had a slope of -1 estudier

hero (hero):

Estudier, even you have to admit it

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

the line I'm looking for has a slope of -1 and is tangent to the circle

OpenStudy (anonymous):

There are 2 such lines...

hero (hero):

http://www.twiddla.com/solved

OpenStudy (anonymous):

3rt2-x = -1?

OpenStudy (anonymous):

The line y = -x...does not have a slope of -1?

OpenStudy (anonymous):

Since when?

OpenStudy (anonymous):

no such line 0.o?

OpenStudy (anonymous):

oh